OpenStudy (anonymous):

P(-4,2) L: 4x+5y=2 What is the equation of a line through P and parallel to L? So far I have y=-4/5x-1(1/5), it says my answer is correct but i need to simplify and use integers or fractions for any numbers in the expression, help please?

OpenStudy (ghazi):

through point (-4,2) and parallel to L so a line parallel to L will be 4x+5y= D now to find D put x=-4, y=2 4*(-4)+5*2=D=-16+10=-6 and your equation is 4x+5y=-6

OpenStudy (anonymous):

thank you sooo much!

OpenStudy (ghazi):

:) YW

OpenStudy (anonymous):

could you help me with one more thing?

OpenStudy (ghazi):

sure

OpenStudy (anonymous):

the follow up question asks for the equation of a line perpendicular instead of parallel, could you explain how to do so?

OpenStudy (anonymous):

it says "what is the equation of a line through P and perpendicular to L"

OpenStudy (ghazi):

for a perpendicular line your slope will be negative off inverse of given line now to find that 4x+5y=2 5y=-4x+2 y=(-4/5)x+(2/5) now your required line is y=(-4/5)x+D to find D put x=-4 and y=2 D= 4/5* -4 +2= -16/5+2= -6/5 y=(-4/5)x+(-6/5)

OpenStudy (anonymous):

for some reason it's telling me that answer is incorrect..

OpenStudy (ghazi):

wait a sec that was wrong...

OpenStudy (ghazi):

slope of your required line is 5/4 and your line is y=(5/4)x+D now put x=-4, y=2 2= 5/4* 4 +D D= 2-5=-3 and your line is y=(5/4)x-3 or 4y=5x-12

OpenStudy (ghazi):

@taranicolee clear?

OpenStudy (anonymous):

oh wait yeah i caught what happened haha i was just coming back to tell you that i figured it out. thank you so much for your help!! (:

OpenStudy (ghazi):

:) YW

OpenStudy (anonymous):

the final answer will actually be y=5/4x+7